OpenStudy (anonymous):

Can someone explain more about input/output tables? Like how to do equations and stuff? Those confuse me.

OpenStudy (amistre64):

input a number and record it; then solve to see what the output value is, and record it

OpenStudy (amistre64):



OpenStudy (amistre64):

y = f(x) in this case :)

OpenStudy (anonymous):

How do you make an equataion for these? I'm soooo confused.

OpenStudy (amistre64):

helps if you know something about recursions

OpenStudy (amistre64):

but your just determining a relation between input and output

OpenStudy (amistre64):

#20 the output increases by 1 when the input increases by 1 y = x + C; where C is some constant that makes the rest of it work out.; in this case when x=0, y = 2.2 2.2 = 0 + C ; C = 2.2 the relation then is: y = x + 2.2

OpenStudy (amistre64):

#22 can be satisfied with t^2 for inputs

OpenStudy (amistre64):

#21 needs more thought since they play nice in it

OpenStudy (anonymous):

Thanks!! :) This helped alot.